Your roommate is an aspiring novelist and asks your opinion on a matter of physics. The novel's central character is kept awake at night by a leaky faucet. The sink is 20.6cm below the faucet. At the instant one drop leaves the faucet, another strikes the sink below and two more in between on the way down.

How many drops per second are keeping the protagonist awake?
Answer #1

Here is what I solved before, please modify the figures as per your question. Please let me know if you have further questions. Ifthis helps then kindly rate 5-stars.

QuestionDetails:

Your roommate is an aspiring novelist. Because you are takingphysics your roommate asks your opinion on a matter of physics. Itseems the central character in the novel is kept awake at night bya leaky faucet. The sink is 19.6 cInm below the leaky faucet. At the instant one drop leavesthe faucet, another strikes the sink below and two more drops arein between on the way down.

How many drops per second are keeping the annoyed protagonistawake?

     The equation for the falling drop is:

gt . The distance fallen is:

4.91 cm 100cm

12 =-= .04s2 4.9

. Start from the point in time where onedrop hits the sink . It takes .2s for a drop to hit the sink.

Assuming the rate of drop formation is constant , there arethus three drops for every 2     ; or , the rate is:

3drops .2s15 drops .
Answer #2

time required for one drop to fall in sink can be detrmined by

s=ut+1/2at2

since u=0 and a=g

s=1/2gt2

t=0.20Sec

0.20sec required for one drop to travel 20.6cm mean while 3more drops are formed

Drops/sec=3/0.20=15 drops